OpenStudy (anonymous):

does the sum of e^(-n^2) from 0 to infinity converge or diverge? how do you use the direct comparison test?

terenzreignz (terenzreignz):

\[\huge \sum_{n=0}^{\infty}e^{-n^2}=\sum_{n=0}^{\infty}\frac{1}{e^{n^2}}\]

terenzreignz (terenzreignz):

First of all, what do you suspect? Having a vague idea can be useful... Do you suspect that this converges or diverges?

terenzreignz (terenzreignz):

And while you're thinking about that, allow me to refresh you on direct comparison test... suppose you have two non-negative series \[\huge \sum_{n=k}^{\infty}a_n \ \ \ \ \ \text{and} \ \ \ \ \ \ \sum_{n=k}^{\infty}b_n\]Such that there is an N>0 where for all n>N \[\huge a_n \le b_n\] Then if \[\huge \sum_{n=k}^{\infty}a_n\]is divergent, then so is \[\huge \sum_{n=k}^{\infty}b_n\] And if \[\huge \sum_{n=k}^{\infty}b_n\]is convergent, then so is \[\huge \sum_{n=k}^{\infty}a_n\]

OpenStudy (anonymous):

it seems like it should be convergent because n is to 2 which is greater than 1

terenzreignz (terenzreignz):

Stop at "It should be convergent" LOL You're asked to do a direct comparison test, right? What are the convergent series that you know of that have its terms greater than or equal to the terms of this series?

OpenStudy (anonymous):

im not sure that is the problem

terenzreignz (terenzreignz):

To use the Direct Comparison test, if you suspect a series is convergent, try looking for a known convergent series having terms greater than or equal to your original series. If, on the other hand, you suspect a series is divergent, try looking for a known divergent series having terms less than or equal to your original series.

OpenStudy (anonymous):

would the series 1/n^2 work?

terenzreignz (terenzreignz):

Let's see :) Is it true that \[\huge \frac1{e^{n^2}}\le\frac1{n^2}\]?

OpenStudy (anonymous):

yes that is true

terenzreignz (terenzreignz):

You know, of course, that \[\Large e^{n^2}\]increases exponentially, and \[\Large n^2\]is just a polynomial, which doesn't increase nearly as quickly. So... in fact, for n = 1, onwards \[\huge n^2<e^{n^2}\]

terenzreignz (terenzreignz):

So, dividing both sides by n squared... and since n squared is positive \[\huge 1 < \frac{e^{n^2}}{n^2}\]and now dividing both sides by e to the n squared \[\huge \frac1{e^{n^2}}<\frac1{n^2}\]

OpenStudy (anonymous):

but since it is 1/e^n^2 and 1/n^2 that would make 1/(e^(n^2)) less than 1/(n^2) right?

terenzreignz (terenzreignz):

That's right :) And so... by the comparison test...? ^.^

OpenStudy (anonymous):

e^(-n^2) is convergent?

terenzreignz (terenzreignz):

Yeap. That was fun :D

OpenStudy (anonymous):

thank you:)